# Induced three-neutron interactions with low cutoffs for dilute neutron matter

### v1

@@ III.3 Cutoff dependence of the neutron-matter EoS
In Fig. 7(a), we show the EoS as a function of the Fermi momentum for different scale factors $f=\lambda/k_{F}$. Looking at the 2N results (dashed lines), we note that they present some cutoff dependence (i.e., dependence on the scale factor $f$). The cutoff dependence gets strongly reduced when we add the HFB contribution of the SRG induced 3N interactions to the SRG 2N results (solid lines). This confirms that the cutoff dependence of the 2N calculation stems mostly from the missing induced 3N interaction and not from missing higher orders of BMBPT, which is plausible because the third-order BMBPT correction is already very small in the chosen range of scale factors Palaniappan et al. (2023).
In order to understand the cutoff dependence better, we plot in Fig. 7(b) the EoS as a function of the scale factor $f=\lambda/k_{F}$ for different Fermi momenta. We see that within the 2N calculation, the energy strongly decreases with decreasing scale factor (short-dashed lines). Once we include the 3N contribution $\varepsilon_{3\text{HFB}}$ (solid line), this cutoff dependence is strongly reduced. Looking in detail, one sees that the correction computed in HFB, $\varepsilon_{3\text{HFB}}$, reduces the cutoff dependence slightly better than the one computed in HF, $\varepsilon_{3\text{HF}}$, (long-dashed lines). The residual cutoff dependence of the EoS gets weaker as the Fermi momentum increases from $k_{F}=0.2\,\text{fm}{}^{-1}$ to $k_{F}=0.8\,\text{fm}{}^{-1}$. However, notice that for $k_{F}=0.8\,\text{fm}{}^{-1}$ and $\lambda\gtrsim 2\,\text{fm}{}^{-1}$, the matrix elements approach the initial ones, and $\lambda$ no longer reflects the true cutoff which can never become larger than the cutoff $2.5\,\text{fm}{}^{-1}$ of the initial interaction. For $\lambda/k_{F}\lesssim 1.3$, one may have to treat the induced 3N interaction in a more complete way than first-order perturbation theory, e.g., by including it into the self-consistent HFB calculation and into the BMBPT. In principle, it might be also necessary to include induced four-body interactions.
@@
This can be also seen in Fig. 8. Since the residual cutoff dependence can be interpreted as a measure for the accuracy of the many-body calculation, we plot the results as bands that are obtained by varying the scale factor in a reasonable range ($1.3\leq\lambda/k_{F}\leq 2.5$). One can see that, except at very small values of $k_{F}$, the red band corresponding to the new 2N+3N SRG results is clearly narrower than the blue hatched band corresponding to the previous 2N $V_{\text{low-}k}$ results. Apart from that, the 2N + induced 3N SRG and the 2N $V_{\text{low-}k}$ results are in excellent agreement, although they are based on different initial 2N interactions (chiral N4LO in the case of SRG, AV18 in the case of $V_{\text{low-}k}$). However, one should keep in mind that the error band obtained from the cutoff dependence does not take into account the uncertainties inherent to the initial interaction and due to missing physics, in particular, the contribution of the bare 3N interaction.
@@ IV Summary and Outlook
In this work, we obtained the induced 3N interaction in the momentum space HPW basis using the SRG formalism. Since $s$-wave (${}^{1}S_{0}$) is the dominant channel in dilute neutron matter, we limit ourselves to the 3N interaction induced by that channel. We note that the convergence of the HPW expansion of the induced 3N matrix elements at low momenta in the SRG flow is improved if we avoid the intermediate summation in the $VV$ term by exploiting the spectator delta functions.
@@
When the SRG-evolved 2N interaction with density-dependent cutoff (i.e., cutoff scaled with the Fermi momentum) is used as input, the EoS of dilute neutron matter within BMBPT shows a strong cutoff dependence when the cutoff gets close to $k_{F}$. Once the induced 3N contribution is included, we find that this cutoff dependence reduces significantly.
Since the induced 3N interaction turns out to be important at low densities, it would be better to track its effect beyond first-order perturbation theory. Its contribution to the EoS and the singlet pairing gap has to be properly studied by including it self-consistently into the HFB and also into the BMBPT calculation. At higher densities, in addition to the 3N interactions induced by higher 2N partial waves, one needs to also include the bare 3N interaction Hebeler (2021); Lovato et al. (2022).
@@
Finally, let us mention that the approach of low-momentum interaction was also applied to systems of superfluid ultracold fermionic atoms Urban and Ramanan (2021). There, the induced 3N interaction was neglected and it would be interesting to include it and see if it helps to reduce the residual cutoff dependence as it does in the case of neutron matter.

### v2

@@ III.3 Cutoff dependence of the neutron-matter EoS
In Fig. 8(a), we show the EoS as a function of the Fermi momentum for different scale factors $f=\lambda/k_{\text{F}}$. Looking at the 2N results (dashed lines), we note that they present some cutoff dependence (i.e., dependence on the scale factor $f$). The cutoff dependence gets strongly reduced when we add the HFB contribution of the SRG induced 3N interactions to the SRG 2N results (solid lines). This confirms that the cutoff dependence of the 2N calculation stems mostly from the missing induced 3N interaction and not from missing higher orders of BMBPT, which is plausible because the third-order BMBPT correction is already very small in the chosen range of scale factors Palaniappan et al. (2023).
In order to understand the cutoff dependence better, we plot in Fig. 8(b) the EoS as a function of the scale factor $f=\lambda/k_{\text{F}}$ for different Fermi momenta. We see that within the 2N calculation, the energy strongly decreases with decreasing scale factor (short-dashed lines). Once we include the 3N contribution $\varepsilon_{3\text{HFB}}$ (solid line), this cutoff dependence is significantly reduced. Looking in detail, one sees that the correction computed in HFB, $\varepsilon_{3\text{HFB}}$, reduces the cutoff dependence slightly better than the one computed in HF, $\varepsilon_{3\text{HF}}$, (long-dashed lines). The residual cutoff dependence of the EoS gets weaker as the Fermi momentum increases from $k_{\text{F}}=0.2\,\text{fm}{}^{-1}$ to $k_{\text{F}}=0.8\,\text{fm}{}^{-1}$. However, notice that for $k_{\text{F}}=0.8\,\text{fm}{}^{-1}$ and $\lambda\gtrsim 2\,\text{fm}{}^{-1}$, the matrix elements approach the initial ones, and $\lambda$ no longer reflects the true cutoff which can never become larger than the cutoff $2.5\,\text{fm}{}^{-1}$ of the initial interaction. For $\lambda/k_{\text{F}}\lesssim 1.3$, one may have to treat the induced 3N interaction in a more complete way than first-order perturbation theory, e.g., by including it into the self-consistent HFB calculation and into the BMBPT. In principle, it might be also necessary to include induced four-body interactions.
@@
This can be also seen in Fig. 9. Since the residual cutoff dependence can be interpreted as a measure for the accuracy of the many-body calculation, we plot the results as bands that are obtained by varying the scale factor in a reasonable range ($1.3\leq\lambda/k_{\text{F}}\leq 2.5$). One can see that, except at very small values of $k_{\text{F}}$, the red band corresponding to the new 2N+3N SRG results is clearly narrower than the blue hatched band corresponding to the previous 2N $V_{\text{low-}k}$ results. Apart from that, the 2N + induced 3N SRG and the 2N $V_{\text{low-}k}$ results are in excellent agreement, although they are based on different initial 2N interactions (chiral N4LO in the case of SRG, AV18 in the case of $V_{\text{low-}k}$). However, one should keep in mind that the error band obtained from the cutoff dependence does not take into account the uncertainties inherent to the initial interaction and due to missing physics, in particular, the contribution of the bare 3N interaction.
@@ IV Summary and Outlook
The present work is an extension of our preceding work Palaniappan et al. (2023) where we studied dilute neutron matter in the framework of BMBPT with $V_{\text{low-}k}$ interactions evolved to very low cutoffs proportional to $k_{\text{F}}$. This method immediately gives rise to the question whether this approach can be trusted if the induced 3N interactions are neglected. Hence, it was necessary to compute the induced 3N interactions, for which the SRG offers a natural framework. Our main result is that induced 3N interactions are indeed relevant for cutoffs less than $\sim 2k_{\text{F}}$ and that their inclusion reduces significantly the cutoff dependence. Furthermore, the final results for the equation of state (using SRG based on chiral N4LO interactions) agree very well with those obtained in our preceding work using $V_{\text{low-}k}$ based on AV18, which shows the universality of the low-momentum effective interactions.
To achieve this, we computed the induced 3N interaction in the momentum space HPW basis using the SRG formalism. Since $s$-wave (${}^{1}S_{0}$) is the dominant channel in dilute neutron matter, we limit ourselves to the 3N interaction induced by that channel. We note that the convergence of the HPW expansion of the induced 3N matrix elements at low momenta in the SRG flow is improved if we avoid the intermediate summation in the $VV$ term by exploiting the spectator delta functions.
@@
When the SRG-evolved 2N interaction with density-dependent cutoff (i.e., cutoff scaled with the Fermi momentum) is used as input, the EoS of dilute neutron matter within BMBPT shows a strong cutoff dependence when the cutoff gets close to $k_{\text{F}}$. Once the induced 3N contribution is included, we find that this cutoff dependence reduces significantly.
Since the induced 3N interaction turns out to be important at low densities, it would be better to track its effect beyond first-order perturbation theory. Its contribution to the EoS has to be properly studied by including it self-consistently into the HFB and also into the BMBPT calculation. At higher densities, in addition to the 3N interactions induced by higher 2N partial waves, one needs to also include the bare 3N interaction Hebeler (2021); Lovato et al. (2022). We plan to address this problem in a future study.
Further, in addition to the EoS, the pairing gap is an important quantity for astrophysical purposes since it affects the thermal evolution and glitches of neutron stars. Therefore, it is important to address the problem of the pairing gap beyond the HFB level. We are currently working on a diagrammatic method to compute BMBPT corrections to the gap as the anomalous self-energy in the Nambu-Gor’kov formalism Schrieffer (1964). Screening and other corrections are then automatically included at higher orders. As for the EoS, it will be necessary to include not only the induced but also the bare 3N interactions into these calculations.
@@
Finally, let us mention that the approach of low-momentum interaction was also applied to systems of superfluid ultracold fermionic atoms Urban and Ramanan (2021). There, the induced 3N interaction was neglected. It would be interesting to include it and see if it helps to reduce the residual cutoff dependence as it does in the case of neutron matter. Work in this direction is in progress.